Not homework - just curious. If a train is going 50 mph and it takes 17 minutes to pass in front of my at the crossing, how long is the train?
Answer by Alan3354(69443)

If a train is going 50 mph and it takes 17 minutes to pass in front of my at the crossing, how long is the train?
----------
50 mi/hr * 17 min * 1/60 hr/min = 85/6 miles = 14.167 miles
A long train